Be sure to answer all parts. Calculate the molecul
Answer #1

Calcium Carbonate CaCO3

atomic weight of Ca = 40.078

atomic weight of C = 12.0107

atomic weight of three O = 3 x 15.9994 = 47.9982

Molar mass of CaCO3 = 40.078 + 12.0107 + 47.9982 = 100.0869 amu (atomic mass units)

Bromous acid HBrO2 :

atomic weight of H = 1.00794

atomic weight of Br = 79.904

atomic weight of two O = 2 x 15.9994 = 31.9988

Molar mass of HBrO2 = 1.00794 + 79.904 + 31.9988 = 112.903594 amu (atomic mass units)

Caffeine (C8H10N4O2)

atomic weight of eight C = 8 x 12.0107 = 96.0856

atomic weight of ten H = 10 x 1.00794 = 10.0794

atomic weight of four N = 4 x 14.0067 = 56.0268

atomic weight of two O = 2 x 15.9994 = 31.9988

Molar mass of CaCO3 = 96.0856 + 10.0794 + 56.0268 + 31.9988 =194.1906 amu (atomic mass units)